Binomial Tree Farm’s financing includes $6.8 million of bank loans. Its common equity is shown in Binomial’s Annual Report at $6.85 million. It has 470,000 shares of common stock outstanding, which trade on the Wichita Stock Exchange at $16.20 per share. What debt ratio should Binomial use to calculate its WACC or asset beta?. (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

Debt ratio = loan/(shares*price+loan)

=6800000/(470000*16.2+6800000) = 47.18%

phpf6hFBT.png